# Betelgeuse’s Buddy: X-Ray Constraints on the Nature of Ori B

### v1

@@ 1 Introduction
Several identities for the companion fit within the plausible mass range of 0.5–2 M⊙, including a compact objects – specifically either a neutron star (NS) or a white dwarf (WD) – or a standard stellar object. In the latter case, given Betelgeuse’s approximate age of 10 Myr (Joyce et al., 2020) and assuming a co-eval formation channel, the companion would be a pre-main sequence young stellar object (YSO), except in the marginal case of a 2M⊙ companion – though we note that uncertainties in the lengths of YSO tracks are large (Tognelli et al., 2011; Baraffe et al., 2015; Haemmerlé et al., 2019; Amard et al., 2019).
@@ 2.2 New Observations with Chandra
We carried out Chandra Director’s Discretionary Time (DDT) observations (PI: A. O’Grady; Program: 25209014) of Betelgeuse between 2024-12-09 and 2024-12-15 using the High Resolution Camera (HRC-I) for a total of 41.85 ks. The observations were broken into seven different epochs, see Table 2 for details. The Chandra data were retrieved from the Chandra Data Archive (CDA)555https://cda.harvard.edu/chaser/.We re-processed the data using the CIAO v4.17.0 data reduction package with CALDB v4.11.6 (see Fruscione et al., 2006). We first ran the chandra_repro command on each individual epoch to remove bad pixels and de-streak the data. We searched each epoch for periods of background solar flares, but found none. We then performed source detection with wavdetect, finding three sources clearly detected in each epoch, albeit at large off-axis angles of 6′. Due to their large off-axis angles, leading to an extended point spread function (PSF) and poorly localized centroid (1-2″ in both RA and DEC at 68% confidence), we cannot accurately correct the native astrometry of our observations. The localization of each source as computed with wavdetect is consistent with slightly higher quality localizations in the Chandra Source Catalog (CSC) version 2.1 (CSC2.1; Evans et al., 2024). The offset is consistent with the expectations for the absolute astrometric error of HRC-I666https://cxc.cfa.harvard.edu/cal/ASPECT/celmon/ which is accurate to 0.85″ at the 90% confidence level (CL).
@@
To combine the seven epochs, we ran the merge_obs command on the full keV HRC-I energy range. We display the combined image, as well as the individual epochs, in Figure 2. There is no obvious source detected at the proper–motion–corrected location of Betelgeuse (e.g., Harper et al., 2017). In order to quantify this, we define a circular source region of radius 1.5″ (93% encircled energy fraction at 1.5 keV), and an annular background region with inner radius 30″ and outer radius of 60″. The encircled energy fraction was determined using the mkpsfmap tool, which also finds a 90% encircled energy radius of 0.95″, computed at 1.5 keV, which is the effective energy of the Chandra/HRC-I “wide” band. We adopt the larger 1.5″ radius given the likely astrometric error is the same order as the 90% radius. In the 1.5″ source aperture, we find a total of 8 total counts in the source region with 4.9 expected background counts. Using Kraft et al. (1991), we estimate a upper limit of 14.6 source counts, resulting in a count rate of cts s-1 ( keV), where we have applied a standard aperture correction. Due to the low background, this method agrees with the results of the aplimits tool (see, e.g., Kashyap et al., 2010). We note that our choice of source aperture does not have a strong impact on our results, and using a 0.95″ radius source region leads to limits that are deeper by only 20% (e.g., cts s-1 in keV).
@@
We further filtered the energy range of the individual event files for each epoch in order to remove background photons777https://cxc.cfa.harvard.edu/ciao/threads/hrci_bg_spectra/. We selected the PI range , corresponding to approximately keV. Summing the source and background counts between the seven epochs, we find 7 total counts in the source region with 3.9 background counts. The corresponding upper limit on the count rate is cts s-1 ( keV).
@@ 3.5 A Young Stellar Object
Further, the Chandra Orion Ultradeep Project (COUP; Flaccomio et al., 2003a, b; Getman et al., 2005), took extremely deep observations of star forming regions in Orion using CXO, surveyed over 1600 YSOs and characterized their X-ray emission. The targets surveyed ranged in ages from Myr. In an analysis of lightly absorbed and well-characterized YSOs, Preibisch et al. (2005) found a positive trend in X-ray luminosity with mass, but a large scatter (see Figure 3 of Preibisch et al. 2005). This scatter is attributed to the influence of accretion, as accreting YSOs do not follow a luminosity-mass relationship as strongly as non-accreting YSOs. Accreting YSOs also have, in general, slightly lower X-ray luminosities than non-accreting YSOs. We note that is likely not accreting because it likely no longer has a disk. Circumstellar disks can be photoevaporated by UV radiation from nearby OB stars (Johnstone et al., 1998; Adams et al., 2004; Andrews, 2020), and Betelgeuse was a hot O-star for it’s MS lifetime. Also, Kraus et al. (2012) finds that for YSO binary systems with close separations (r40 AU – the separation of is only 9AU), 2/3rds of these systems disperse their disks within 1 Myr of formation.
@@ 4 Conclusions
We have presented our analysis of recent Chandra observations of Betelgeuse in December 2024 during the period of maximum separation between Betelgeuse and its proposed companion. We do not detect an X-ray source at the proper motion corrected position of Betelgeuse and instead place upper limits on the underlying X-ray luminosity of roughly erg s-1 ( to ) in keV for the expected hydrogen column densities cm-2 from Betelgeuse’s wind. This limit is robust to different model assumptions under this specific value. These X-ray limits exclude a compact object (e.g., white dwarf, neutron star) as the companion to Betelgeuse. While we therefore strongly favor a young stellar object as the companion, we cannot place strong constraints on its mass, though a sub-solar mass companion is slightly favored (see Goldberg et al. 2025 for further discussion of complementary UV constraints from the Hubble Space Telescope).
@@ Acknowledgements
The authors thank Andrea Dupree, Saavik Ford, Yuri Levin, Morgan MacLeod, Barry McKernan, Koji Mukai, Adiv Paradise, Mathieu Renzo, and Leike van Son for useful discussions. A.O. also acknowledges the support of Jupiter O’Grady.
@@
This research made use of NASA’s Astrophysics Data System Bibliographic Services, the pathfinder framework (Iyer et al., 2024) as well as of the SIMBAD database and the cross-match service operated at CDS, Strasbourg, France.

### v2

@@ 1 Introduction
Several identities for the companion fit within the plausible mass range of 0.5–2 M⊙, including a compact objects – specifically either a neutron star (NS) or a white dwarf (WD) – or a standard stellar object. We note that smaller objects, such as brown dwarfs or large Jupiter-class planets, are ruled out by the MacLeod et al. (2025); Goldberg et al. (2024) mass constraints. In the stellar object case, given Betelgeuse’s approximate age of 10 Myr (Joyce et al., 2020) and assuming a co-eval formation channel, the companion would be a pre-main sequence young stellar object (YSO), except in the marginal case of a 2M⊙ companion – though we note that uncertainties in the lengths of YSO tracks are large (Tognelli et al., 2011; Baraffe et al., 2015; Haemmerlé et al., 2019; Amard et al., 2019).
@@ 2.2 New Observations with Chandra
We carried out Chandra Director’s Discretionary Time (DDT) observations (PI: A. O’Grady; Program: 25209014555https://doi.org/10.25574/cdc.424) of Betelgeuse between 2024-12-09 and 2024-12-15 using the High Resolution Camera (HRC-I) for a total of 41.85 ks. The observations were broken into seven different epochs, see Table 2 for details. The Chandra data were retrieved from the Chandra Data Archive (CDA)666https://cda.harvard.edu/chaser/.We re-processed the data using the CIAO v4.17.0 data reduction package with CALDB v4.11.6 (see Fruscione et al., 2006). We first ran the chandra_repro command on each individual epoch to remove bad pixels and de-streak the data. We searched each epoch for periods of background solar flares, but found none. We then performed source detection with wavdetect, finding three sources clearly detected in each epoch, albeit at large off-axis angles of 6′. Due to their large off-axis angles, leading to an extended point spread function (PSF) and poorly localized centroid (1-2″ in both RA and DEC at 68% confidence), we cannot accurately correct the native astrometry of our observations. The localization of each source as computed with wavdetect is consistent with slightly higher quality localizations in the Chandra Source Catalog (CSC) version 2.1 (CSC2.1; Evans et al., 2024). The offset is consistent with the expectations for the absolute astrometric error of HRC-I777https://cxc.cfa.harvard.edu/cal/ASPECT/celmon/ which is accurate to 0.85″ at the 90% confidence level (CL).
@@
To combine the seven epochs, we ran the merge_obs command on the full keV HRC-I energy range. We display the combined image, as well as the individual epochs, in Figure 2. There is no obvious source detected at the proper–motion–corrected location of Betelgeuse (e.g., Harper et al., 2017). In order to quantify this, we define a circular source region of radius 1.5″ (93% encircled energy fraction at 1.5 keV), and an annular background region with inner radius 30″ and outer radius of 60″. The encircled energy fraction was determined using the mkpsfmap tool, which also finds a 90% encircled energy radius of 0.95″, computed at 1.5 keV, which is the effective energy of the Chandra/HRC-I “wide” band. We adopt the larger 1.5″ radius given the likely astrometric error is the same order as the 90% radius. In the 1.5″ source aperture, we find a total of 8 total counts in the source region with 4.9 expected background counts. Using the aplimits tool (Kashyap et al., 2010), we estimate a count rate upper limit of cts s-1 ( keV), where we have applied a standard aperture correction. We note that our choice of source aperture does not have a strong impact on our results, and using a 0.95″ radius source region leads to limits that are deeper by only 20% (e.g., cts s-1 in keV).
@@
We further filtered the energy range of the individual event files for each epoch in order to remove background photons888https://cxc.cfa.harvard.edu/ciao/threads/hrci_bg_spectra/. We selected the PI range , corresponding to approximately keV. Summing the source and background counts between the seven epochs, we find 7 total counts in the source region with 3.9 background counts. The corresponding upper limit on the count rate is cts s-1 ( keV).
@@ 3.1 X-ray Absorption from the Surrounding Environment
As a final check, we consider the standard relation between optical extinction and hydrogen column density in our Galaxy (Predehl & Schmitt, 1995; Güver & Özel, 2009), applying most generally to the interstellar medium. Using the visual extinction ( mag; Montargès et al. 2021) measured for Betelgeuse (containing contributions from the interstellar medium out to 168 pc and the contribution from its surrounding environment and wind), we find an estimate of cm-2. It is not a given that the surrounding environment and wind of Betelgeuse lie on this relation as the total Hydrogen column in the star’s circumstellar environment likely contributes to X-ray absorption in addition to the dust that contributes to the measured visual extinction. However, if the standard relation holds, it would show that most of our estimates, as expected, are overconservative and that our conclusions should become more robust. Based on Figure 3, for the majority of models this smaller value would decrease the upper limits be a factor of a few.
@@ 3.5 A Young Stellar Object
Further, the Chandra Orion Ultradeep Project (COUP; Flaccomio et al., 2003a, b; Getman et al., 2005), took extremely deep observations of star forming regions in Orion using Chandra, surveyed over 1600 YSOs and characterized their X-ray emission. The targets surveyed ranged in ages from Myr. In an analysis of lightly absorbed and well-characterized YSOs, Preibisch et al. (2005) found a positive trend in X-ray luminosity with mass, but a large scatter (see Figure 3 of Preibisch et al. 2005). This scatter is attributed to the influence of accretion, as accreting YSOs do not follow a luminosity-mass relationship as strongly as non-accreting YSOs. Accreting YSOs also have, in general, slightly lower X-ray luminosities than non-accreting YSOs. We note that is likely not accreting because it likely no longer has a disk. Circumstellar disks can be photoevaporated by UV radiation from nearby OB stars (Johnstone et al., 1998; Adams et al., 2004; Andrews, 2020), and Betelgeuse was a hot O-star for it’s MS lifetime. Also, Kraus et al. (2012) finds that for YSO binary systems with close separations (r40 AU – the separation of is only 9AU), 2/3rds of these systems disperse their disks within 1 Myr of formation.
@@ 4 Conclusions
We have presented our analysis of recent Chandra observations of Betelgeuse in December 2024 during the period of maximum separation between Betelgeuse and its proposed companion. We do not detect an X-ray source at the proper motion corrected position of Betelgeuse and instead place upper limits on the underlying X-ray luminosity of roughly erg s-1 ( to ) in keV for the expected hydrogen column densities cm-2 from Betelgeuse’s wind. This limit is robust to different model assumptions under this specific value. These X-ray limits exclude a compact object (e.g., white dwarf, neutron star) as the companion to Betelgeuse. While we therefore strongly favor a young stellar object as the companion, we cannot place strong constraints on its mass, though a sub-solar mass companion is slightly favored (see Goldberg et al. 2025 for further discussion of complementary UV constraints from the Hubble Space Telescope). Finally, we note that during the review process for this article, a possible direct-image detection of the companion was published (Howell et al., 2025). This work suggests that the companion is a stellar object with a mass of 1.4-2M⊙, which is within the overall predicted mass range, and specifically consistent with our results that the companion cannot be a compact object.
@@ Acknowledgements
The authors thank Andrea Dupree, Saavik Ford, Yuri Levin, Morgan MacLeod, Barry McKernan, Koji Mukai, Adiv Paradise, Mathieu Renzo, and Leike van Son for useful discussions. A.O. also acknowledges the support of Jupiter O’Grady. The authors also thank the anonymous reviewer for a helpful and constructive referee report.
@@
This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection https://doi.org/10.25574/cdc.424 (catalog doi:10.25574/cdc.424). The scientific results reported in this article are based on observations made by the Chandra X-ray Observatory (CXO). This research has made use of data obtained from the Chandra Data Archive provided by the Chandra X-ray Center (CXC). This research has made use of software provided by the Chandra X-ray Center (CXC) in the application package CIAO. This research made use of NASA’s Astrophysics Data System Bibliographic Services, the pathfinder framework (Iyer et al., 2024) as well as of the SIMBAD database and the cross-match service operated at CDS, Strasbourg, France.